10        20        30        40        50        60        70 T0076 1wdcC  STDDSPYKQAFSLFDRHGTGRIPKTSIGDLLRACGQNPTLAEITEIESTLPAE----------------- ----------------------LS------------------QDEIDDLKDVFELFDFWDGRDGAVDAFK 80        90       100       110       120       130     140 T0076 1wdcC  --------------------------------VDMEQFLQVLNRPNGFDMPGDPEEFVKGFQVFDK---D LGDVCRCLGINPRNEDVFAVGGTHKMGEKSLPFEEFLPAYEGLM------DCEQGTFADYMEAFKTFDRE 150       160       170       180       190 T0076 1wdcC  ATGMIGVGELRYVLTS---LGEKLSNEEMDELLKGVPV--KDGMVNYHDFV--QMILAN GQGFISGAELRHVLTALGERLSDEDVDEIIKLTDLQEDLEGNVKYEDFVKKVMAGPYP-

Figure 1: Comparison of candidate 1wdcC-t76-const-global.pw.a2m.gz (lines) to reference t76-1wdcC-GL.a2m (letters). Reference alignment has 93 matching columns, Candidate alignment has 133 matching columns. 173 were used (for template shift+target shift) to compute shift score. The shift score is 0.219801, with a min shift of 0, max shift of 47, and mean of 6.046707. Optimal subalignment has 61 matching columns, with 41 paired matches. 82 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.338060, with a min shift of 0, max shift of 3, and mean of 1.073171. Optimum achieved by removing the 92 columns indicated with dotted lines from the candidate alignment.